OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 40°C and 180°C. The temperature is at its lowest point when t = 0 and completes one cycle over a 12-hour period. What is a sine function that would model this reaction? f(x) = 70 sin 12x + 110 f(x) = 110 sin 12x + 70 f(x) = 70 sin pi over 6x + 110 f(x) = 110 sin pi over 6x + 70

OpenStudy (anonymous):

I said B but that was incorrect. I'm leaning more towards C now.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you need to better notate the options ... use grouping symbols to clarify what it stuff

OpenStudy (anonymous):

\[A. f(x)= 70\sin 12x + 110\]

OpenStudy (anonymous):

\[B. f(x)= 110\sin 12x+ 70\]

OpenStudy (anonymous):

\[C. f(x) = 70 \sin \frac{ \pi }{ 6} +110\]

OpenStudy (anonymous):

\[D. f(x)= 110 \sin \frac{ \pi }{ 6 }x + 70\]

OpenStudy (anonymous):

sorry for the first 3 equations there should be a space between the number and sin

OpenStudy (amistre64):

sin is a function: like f when we write in function notation we say f(this) to define what the function operates on. we can then decipher when multiple functions are at play:\[sin(x)+(2x)^2-tan(\frac{pi}4x+2)\] if this is written in any other way, it becomes vague \[\sin x+2x^2-\tan\frac{pi}4x+2\]

OpenStudy (amistre64):

\[\sin (x+2x^2)-\tan(\frac{pi}4x+2)\]

OpenStudy (anonymous):

ok. Sorry. If It takes a while for me to respond but I am taking notes. :)

OpenStudy (anonymous):

ok. so far so good. I'm a pretty fast writer and I abbreviate a lot :) Plus I can go back later and add in anything I might have missed so you can move at your own normal place

OpenStudy (amistre64):

one complete cycle is a period, agreed?

OpenStudy (amistre64):